OpenStudy (anonymous):

I don't understand how the below function is undefined at 0...

OpenStudy (anonymous):

\(\ \Huge f(x)=sin(\frac{\pi}{x}).\)

OpenStudy (lgbasallote):

uhh because \(\sin (\frac{\pi}{0})\) is undefined?

OpenStudy (anonymous):

@lgbasallote I'm trying to understand \(\ \Huge \lim_{x \rightarrow 0} sin (\frac{\pi}{x}) \neq 0. \)

OpenStudy (lgbasallote):

oh... o.O

OpenStudy (unklerhaukus):